# Read Between the Layers: Leveraging Intra-Layer Representations for Rehearsal-Free Continual Learning with Pre-Trained Models

### v2

@@ 4.2 Combination with parameter-efficient model adaptation
Collect $\Phi_{-k:}({\bm{x}}_{t})$ (Eq. 4)
${\bm{G}}\leftarrow{\bm{G}}+\Phi_{-k:}({\bm{x}}_{t})\otimes\Phi_{-k:}({\bm{x}}_{t})$
${\bm{c}}_{y_{t}}\leftarrow{\bm{c}}_{y_{t}}+\Phi_{-k:}({\bm{x}}_{t})$EndFor
@@
To maintain the powerful generic features of the pre-trained backbone, we choose to apply FSA to an additional set of learnable PETL parameters while keeping the backbone frozen throughout. To update PETL parameters during FSA, we train a linear classification head with $N_{1}$ outputs via Adam optimizer (Kingma & Ba, 2015) and cross-entropy loss only during the first CL stage and discard the classification head afterward. During the CL phase, we keep all model parameters frozen and update ${\bm{G}}$ and ${\bm{c}}_{y}$ only. We follow prior works (Zhou et al., 2023b; McDonnell et al., 2023) and experiment with Visual Prompt Tuning (VPT) (Jia et al., 2022), Scaling and Shifting of Features (SSF) (Lian et al., 2022), and AdaptFormer (Chen et al., 2022) as PETL methods and refer to the cited papers for details.
@@ 5 Experiments
In what follows, we conduct a series of experiments to assess our approach under various CL settings. We start with an overview of the datasets and implementation details in Sec. 5.1, followed by empirical comparisons of our method against recent strong baselines under CIL, DIL, and OCL settings in Sec. 5.2 and Sec. 5.3. Ablation studies are presented in Sec. 5.4. We explore the impact of different configurations of the last $k$ layers on classifier performance in Sec. 5.5 and conclude our empirical analysis by investigating the range of classes that benefit from intra-layer features in Sec. 5.6.
@@ 5.1 Datasets and implementation details
Following prior works (Wang et al., 2022b; c; Zhou et al., 2023b; Zhang et al., 2023; McDonnell et al., 2023), we experiment with two ViT-B/16 (Dosovitskiy et al., 2021) models, the former with self-supervised pre-training on ImageNet-21K (ViT-B/16-IN21K) (Ridnik et al., 2021), the latter with additional supervised fine-tuning on ImageNet-1K (ViT-B/16-IN1K) (Krizhevsky et al., 2012). During adaptation in the CIL and DIL settings, PETL parameters are trained using a batch size of 48 for 20 epochs, Adam (Kingma & Ba, 2015) with momentum for optimization, and learning rate scheduling via cosine annealing, starting with 0.03. We train five prompt tokens for VPT and use a bottleneck dimension of 16 for AdaptFormer. All baselines are rehearsal-free and trained on the same aforementioned ViT backbones.
@@ 5.2 Performance in the CIL and DIL settings
Table 1: Comparison of prompting, backbone fine-tuning, and class-prototype methods for the CIL setting. Results are taken from McDonnell et al. (2023) except results for SLCA marked with (), which are reproduced using the officially released code.
@@
In the CIL setting, we compare LayUP with several prompt learning, fine-tuning, and class-prototype methods for CL. We additionally report results for joint fine-tuning (FT) of the full backbone and a linear probe. As shown in Tab. 1, LayUP surpasses all baselines for four of the seven split datasets. Interestingly, these four datasets (IN-R, IN-A, VTAB, Cars) have two distinctive characteristics: First, they have the highest domain gap among all datasets, as indicated by the cosine distance to the pre-trained miniImageNet domain for the offline setting and 50 shots (cf. Panos et al. (2023)). Second, they have significantly less training data than CIFAR, OB, and CUB (cf. Appendix B). The former confirms our initial hypothesis that intra-layer representations are more domain-invariant and consequently more robust to large distributional shifts from the source to the target domain. The latter indicates that especially in the low-data regime, in contrast to other methods that tend to overfit to the target domain, LayUP constructs class prototypes and decision boundaries that generalize well even if the amount of training data is scarce compared with the data used for ViT pre-training. It is further noteworthy that RanPAC, which is the only baseline that LayUP does not consistently outperform, is considerably more expensive regarding both memory and computation (cf. Sec. 4.3).
@@ 5.3 Performance in the OCL setting
5.4 Ablation study
The purpose of the following ablation study is to demonstrate the advantages of leveraging intra-layer representations over solely utilizing features from the final representation layer (i.e., $k=1$) for class-prototype construction in CL. As LayUP integrates intra-layer representations with second-order statistics via Gram matrix inversion as well as first-session adaptation in the CIL and DIL settings, we consider three configurations for ablating intra-layer representations: (i) in the standard setting with both FSA and second-order statistics (LayUP vs. $k=1$), (ii) excluding the FSA stage, i.e., omitting phase A in Alg. 1 (w/o FSA vs. $k=1$, w/o FSA), and (iii) excluding FSA and second-order statistics, which corresponds to traditional nearest-mean classifiers (LayNMC vs. NMC). Results for the baselines constructed for settings (i)-(iii) can be found in the bottom sections of Tab. 1 and Tab. 2. For (i), enriching the final layer with intermediate features consistently results in absolute performance gains ranging from 1.2% to 5.7% (CIL) and 2.2% to 5.2% (DIL). For (ii), except for the VTAB dataset in the CIL setting, leveraging intra-layer features yields absolute accuracy improvements of up to 9.5% (CIL) and 9.0% (DIL). Lastly, for (iii), LayNMC does not show superior performance to the NMC baseline. Such results indicate that the benefits of intermediate features diminish when not decorrelated via Gram matrix inversion, as they contain more lower-level noisy information than final layer representations.
@@ 5.4 Ablation study
In the lower section of Tab. 3, further comparisons are made between ablated versions of LayUP and prototypes using final representations ($k=1$) in the OCL setting for two selected values of the regularization parameter $\lambda$. LayUP generally outperforms its ablated versions, although it demonstrates decreased performance in low-data benchmarks such as IN-A and VTAB, particularly when omitting regularization ($\lambda=0$). This decline in performance may stem from the increased dimensionality of class prototypes and the Gram matrix due to the addition of intra-layer representations, thereby elevating the risk of overfitting. Consequently, the absence of regularization, combined with limited training data, may render intra-layer features detrimental. However, in any other case, they enhance classification accuracy by 0.6% to 8.3%.
5.5 Analysis of intra-layer representation depth
To provide more insights into the impact of the intra-layer representation depth on LayUP performance, we plot in Tab. 4 the average accuracy after training for different choices of $k$ using a ViT-B/16-IN1K backbone and AdaptFormer as PETL method.
@@ 5.5 Analysis of intra-layer representation depth
The final layer classifier ($k=1$) exhibits the lowest performance in comparison to all other $k$ values across all datasets, except for the split VTAB dataset. Conversely, LayUP configurations with medium or large $k$ values almost consistently outperform those with small $k$ values, demonstrating the broad advantages of leveraging representations from multiple intermediate layers for class-prototype construction. While the optimal choice of $k$ varies across datasets, performance generally improves or stabilizes with increasing $k$ values, thereby selecting intermediate features from a higher number of consecutive layers. However, a larger $k$ value corresponds to increased memory and runtime complexity during inference, necessitating careful consideration of its choice based on the demands and constraints at different dimensions.
5.6 How universal are intra-layer prototypes?
@@
We aim to assess the universality of intra-layer prototypes for classification, specifically, the breadth of classes benefiting from LayUP. To this end, we report the proportion of classes for each dataset whose test accuracy scores are higher, equal, or lower with LayUP compared to the final layer ridge classifier (cf. Eq. 2). Beyond merely counting the classes that are better predicted by either method, we also examine the extent of the benefit one method has over the other per improved class. Consequently, we calculate the average difference in accuracy scores between LayUP and the final layer classifier ($k=1$), and vice versa, for each class that shows improvement.
@@
Figure 6: LayUP vs. classification from the final representation layer. Left: Percentage of classes per dataset that are better classified with LayUP (with $k=6$) compared with the last layer ridge classifier (equivalent to LayUP with $k=1$). Right: Difference of absolute accuracy per improved class between LayUP and the last layer classifier ($k=1$), and vice versa.
@@
Fig. 6 illustrates that a higher percentage of classes (between 18% and 72%) is more effectively classified upon introducing intermediate representations to class-prototype construction. Although the difference in the number of improved classes is not as pronounced for IN-A and CUB compared with all other benchmarks, the relative difference in accuracy per improved class between LayUP and the final representation layer classifier reaches as high as 39% for IN-A and 22% for CUB. This indicates that for these two datasets, the introduction of intra-layer representations significantly benefits a few classes rather than slightly benefiting a large number of classes, as observed with the other benchmarks. Despite the variability with respect to the number of classes per dataset that benefit from LayUP and the extent of the benefit, our results demonstrate the universal applicability of intra-layer prototypes across a broad spectrum of tasks, classes, or domains.
@@ 6 Conclusion
In this paper, we propose LayUP, a simple yet effective rehearsal-free class-prototype method for continual learning with pre-trained models. LayUP leverages intra-layer representations of a pre-trained feature extractor to increase robustness and generalizability, especially under large domain gaps and in low-data regimes. It further computes second-order feature statistics to decorrelate class prototypes, combined with parameter-efficient first-task adaptation. Extensive experiments across a range of image classification datasets and continual learning settings demonstrate that LayUP performs strongly against competitive baselines while requiring significantly less memory and computational complexity. In future work, we will further investigate integrating class-prototype methods with continual adaptation of pre-trained models beyond first-task adaptation, which is particularly beneficial for learning under multiple distributional shifts. Intra-layer representations serve as a potent source of knowledge that is inherently available in pre-trained models. Based on the findings made in this work, we encourage further exploration into making explicit use of these representations (i.e., reading between the layers) in future approaches to continual learning.

### v3

@@ 4.2 Combination with parameter-efficient model adaptation
# Initialization of class prototypes
${\bm{G}}\leftarrow{\bm{0}}\in\mathbb{R}^{d_{(k)}\times d_{(k)}}$
${\bm{c}}_{y}\leftarrow{\bm{0}}\in\mathbb{R}^{d_{(k)}}\ \forall\ y\in\mathcal{Y}$
@@
Collect $\Phi_{-k:}({\bm{x}})$ (Eq. 5)
${\bm{G}}\leftarrow{\bm{G}}+\Phi_{-k:}({\bm{x}})\otimes\Phi_{-k:}({\bm{x}})$
${\bm{c}}_{y}\leftarrow{\bm{c}}_{y}+\Phi_{-k:}({\bm{x}})$EndFor
@@
To maintain the powerful generic features of the pre-trained backbone, we choose to apply FSA to an additional set of learnable PETL parameters while keeping the backbone frozen throughout. To update PETL parameters during FSA, we train a linear classification head with as many output neurons as the number of unique labels in the first task via Adam optimizer (Kingma & Ba, 2015) and cross-entropy loss only during the first CL stage and discard the classification head afterward. During the CL phase, we keep all model parameters frozen and update ${\bm{G}}$ and ${\bm{c}}_{y}$ only. We follow prior works (Zhou et al., 2023b; McDonnell et al., 2023) and experiment with Visual Prompt Tuning (VPT) (Jia et al., 2022), Scaling and Shifting of Features (SSF) (Lian et al., 2022), and AdaptFormer (Chen et al., 2022) as PETL methods and refer to the cited papers for details.
@@ 5 Experiments
In what follows, we conduct a series of experiments to assess our approach under various CL settings. We start with an overview of the datasets and implementation details in Sec. 5.1, followed by empirical comparisons of our method against recent strong baselines under CIL, DIL, and OCL settings in Sec. 5.2 and Sec. 5.3. Ablation studies are presented in Sec. 5.4. We explore the impact of different configurations of the last $k$ layers on classifier performance in Sec. 5.5 and investigate the range of classes that benefit from hidden features in Sec. 5.6. We conclude our empirical analysis by evaluating LayUP as a plug-in for other class-prototype methods in Sec. 5.7.
@@ 5.1 Datasets and implementation details
Following prior works (Wang et al., 2022b; c; Zhou et al., 2023b; Zhang et al., 2023; McDonnell et al., 2023), we experiment with two ViT-B/16 (Dosovitskiy et al., 2021) models, the former (ViT-B/16-IN21K) with self-supervised pre-training on the ImageNet-21K dataset (Ridnik et al., 2021), the latter (ViT-B/16-IN1K) with additional supervised fine-tuning on the ImageNet-1K dataset (Krizhevsky et al., 2012). During adaptation in the CIL and DIL settings, PETL parameters are trained using a batch size of 48 for 20 epochs, Adam (Kingma & Ba, 2015) with momentum for optimization, and learning rate scheduling via cosine annealing, starting with 0.03. We train five prompt tokens for VPT and use a bottleneck dimension of 16 for AdaptFormer. All baselines are rehearsal-free and trained on the same aforementioned ViT backbones.
@@ 5.2 Performance in the CIL and DIL settings
Table 1: Comparison of prompting, backbone fine-tuning, and class-prototype methods for the CIL setting. Results are taken from McDonnell et al. (2023) except results for SLCA marked with (), which are reproduced using the officially released code. Best results per dataset are highlighted in bold.
@@
In the CIL setting, we compare LayUP with several prompt learning, fine-tuning, and class-prototype methods for CL. We additionally report results for joint fine-tuning (FT) of the full backbone and a linear probe. As shown in Tab. 1, LayUP surpasses all baselines for four of the seven split datasets.
The four datasets—IN-R, IN-A, VTAB, and Cars—on which LayUP consistently outperforms all baseline models, exhibit two distinct characteristics: First, they possess a high domain gap relative to the pre-training ImageNet domain, as detailed in Sec. 5.5. Second, with the exception of IN-R, they contain significantly less training data compared to CIFAR, OB, and CUB (cf. Appendix B). The former confirms our initial hypothesis that intermediate representations are more domain-invariant and consequently more robust to large distributional shifts from the source to the target domain. The latter indicates that especially in the low-data regime, in contrast to other methods that tend to overfit to the target domain, LayUP constructs class prototypes and decision boundaries that generalize well even if the amount of training data is scarce compared with the data used for ViT pre-training. It is further noteworthy that RanPAC, which is the only baseline that LayUP does not consistently outperform, is considerably more expensive regarding both memory and computation (cf. Sec. 4.3).
@@ 5.3 Performance in the OCL setting
5.4 Ablation study
The purpose of the following ablation study is to demonstrate the advantages of leveraging multi-layer representations over solely utilizing features from the final representation layer (i.e., $k=1$) for class-prototype construction in CL. As LayUP integrates multi-layer representations with second-order statistics via Gram matrix inversion as well as first session adaptation in the CIL and DIL settings, we consider three configurations for ablating multi-layer representations: (i) in the standard setting with both FSA and second-order statistics (LayUP vs. $k=1$), (ii) excluding the FSA stage, i.e., omitting phase A in Alg. 1 (w/o FSA vs. $k=1$, w/o FSA), and (iii) excluding FSA and second-order statistics, which corresponds to traditional nearest-mean classifiers (LayNMC vs. NMC). Results for the baselines constructed for settings (i)-(iii) can be found in the bottom sections of Tab. 1 and Tab. 2. For (i), enriching the final layer with intermediate features consistently results in absolute performance gains ranging from 1.2% to 5.7% (CIL) and 2.2% to 5.2% (DIL). For (ii), except for the VTAB dataset in the CIL setting, leveraging multi-layer features yields absolute accuracy improvements of up to 9.5% (CIL) and 9.0% (DIL). Lastly, for (iii), LayNMC does not show superior performance to the NMC baseline. Such results indicate that the benefits of intermediate features diminish when not decorrelated via Gram matrix inversion, as they contain more lower-level noisy information than final layer representations.
@@ 5.4 Ablation study
In the lower section of Tab. 3, further comparisons are made between ablated versions of LayUP and prototypes using final representations ($k=1$) in the OCL setting for two selected values of the regularization parameter $\lambda$. LayUP generally outperforms its ablated versions, although it demonstrates decreased performance in low-data benchmarks such as IN-A and VTAB, particularly when omitting regularization ($\lambda=0$). This decline in performance may stem from the increased dimensionality of class prototypes and the Gram matrix due to the addition of intermediate representations, thereby elevating the risk of overfitting. Consequently, the absence of regularization, combined with limited training data, may render multi-layer features detrimental. However, in any other case, they enhance classification accuracy by 0.6% to 8.3%.
5.5 Analysis of multi-layer representation depth
@@
To provide more insights into the impact of the multi-layer representation depth on LayUP performance, we plot in Tab. 4 the average accuracy after training for different choices of $k$ using a ViT-B/16-IN1K backbone and AdaptFormer as PETL method. The final-layer classifier ($k=1$) exhibits the lowest performance in comparison to all other $k$ values across all datasets, except for the split VTAB dataset. Conversely, LayUP configurations with medium or large $k$ values almost consistently outperform those with small $k$ values, demonstrating the broad advantages of leveraging representations from multiple intermediate layers for class-prototype construction. While the optimal choice of $k$ varies across datasets, the performance gains generally diminish with increasing $k$ values. For instance, the average accuracy gain from choosing $k=6$ over $k=1$ is $2.7\%$, whereas it is only $0.4\%$ when choosing $k=12$ over $k=6$. This confirms the findings made in Sec. 4.1, suggesting that a value of $k=6$ might generally suffice to achieve a significant performance gain over last-layer class-prototype methods while requiring significantly less computational and memory resources compared with $k=12$.
We further aim to determine whether prior knowledge about the characteristics of downstream CL data can inform the optimal choice of maximum multi-layer representation depth with respect to overall performance. Specifically, we identify the value of $k$ that maximizes accuracy on the test set after running Alg. 1 with AdaptFormer as PETL method for each split dataset used in Sec. 5.2 and for each pre-trained ViT model. We plot the highest performing $k$ per dataset against three measures: (i) the degree of domain gap to the pre-training ImageNet domain, measured by Maximum Mean Discrepancy (MMD); (ii) the intra-class similarity, measured by the average pairwise cosine similarity between training image feature vectors of the same class; and (iii) the inter-class similarity, measured by the average pairwise cosine similarity between image feature vectors of disjoint classes. For (i), we follow Panos et al. (2023) and use the miniImageNet dataset (60K images) as a proxy for the ImageNet-1K (1.3M images) and ImageNet-21K (14M images) pre-training datasets to reduce computational overhead when calculating pairwise distances. A large MMD value indicates that two datasets have significantly different statistical properties, thus signifying a large domain gap between the source and target domains.
Figure 6: Choice of $k$ for different dataset characteristics. Multi-layer representation depth $k$ that yields highest accuracy vs. normalized MMD between each dataset and the ImageNet pre-training domain, represented by miniImageNet dataset (left), intra-class similarity (center), and inter-class similarity (right).
As shown in Fig. 6, there is no clear relationship between the MMD and the optimal choice of $k$. This suggests that the degree of domain gap from the source domain of a pre-trained model is not a reliable indicator for determining the multi-layer representation depth that maximizes performance. However, a positive trend is observed between the optimal $k$ and both intra-class and inter-class similarity.
First, datasets with high intra-class and inter-class similarity (e.g., Cars and CUB) benefit from a high value of $k$. These datasets typically focus on fine-grained classification of natural images within a specific domain. Second, datasets with medium intra-class similarity and low inter-class similarity (e.g., VTAB, OB, or CIFAR) benefit from a medium-to-high value of $k$. These datasets are usually composed of multiple specialized natural-image datasets, which can belong to very distinct domains, thus explaining the low inter-class similarity. Finally, datasets with low intra-class and inter-class similarity (e.g., IN-A or IN-R) benefit from a medium value of $k$. These datasets often consist of atypical or stylized image examples that differ from the natural images in ImageNet and are thus prone to be distributed unsystematically in the feature space.
Although the maximum performance for most configurations in Fig. 6 is achieved by concatenating features from the majority of layers in the pre-trained model, the performance differences between various choices of $k$ can be minimal in some cases (as shown in Tab. 4) and may fluctuate with different random initializations. Furthermore, as previously discussed, performance gains tend to saturate as $k$ increases. Concurrently, larger $k$ values are associated with higher memory usage and increased runtime complexity during inference. Therefore, selecting an appropriate $k$ requires careful consideration of the specific demands and constraints across different dimensions.
5.6 How universal are multi-layer prototypes?
@@
We aim to assess the universality of multi-layer prototypes for classification, specifically, the breadth of classes benefiting from LayUP. To this end, we report the proportion of classes for each dataset whose test accuracy scores are higher, equal, or lower with LayUP compared to the final-layer ridge classifier (cf. Eq. 3). Beyond merely counting the classes that are better predicted by either method, we also examine the extent of the benefit one method has over the other per improved class. Consequently, we calculate the average difference in accuracy scores between LayUP and the final-layer classifier ($k=1$), and vice versa, for each class that shows improvement.
@@
Figure 7: LayUP vs. classification from the final representation layer. Left: Percentage of classes per dataset that are better classified with LayUP (with $k=6$) compared with the last layer ridge classifier (equivalent to LayUP with $k=1$). Right: Difference of absolute accuracy per improved class between LayUP and the final-layer classifier ($k=1$), and vice versa.
@@
Fig. 7 illustrates that a higher percentage of classes (between 18% and 72%) is more effectively classified upon introducing intermediate representations to class-prototype construction. Although the difference in the number of improved classes is not as pronounced for IN-A and CUB compared with all other benchmarks, the relative difference in accuracy per improved class between LayUP and the final representation layer classifier reaches as high as 39% for IN-A and 22% for CUB. This indicates that for these two datasets, the introduction of intermediate representations significantly benefits a few classes rather than slightly benefiting a large number of classes, as observed with the other benchmarks. Despite the variability with respect to the number of classes per dataset that benefit from LayUP and the extent of the benefit, our results demonstrate the universal applicability of multi-layer prototypes across a broad spectrum of tasks, classes, or domains.
5.7 Combination with other class-prototype methods
[…]
The results for the combinations with ADAM and RanPAC are shown in Tab. 5(a) and Tab. 5(b), respectively. Integrating LayUP with other class-prototype methods consistently improves performance across all datasets. However, the magnitude of these improvements varies, with the CUB and CIFAR datasets showing the least pronounced enhancements ($\uparrow 0.4\%$ for LayUP integrated with RanPAC) and the Cars dataset exhibiting the most substantial gains ($\uparrow 31.1\%$ for LayUP integrated with ADAM). Although the optimal value of $k$ for maximizing performance differs across datasets, the overall performance does not significantly vary with different $k$ values. This suggests that a small, resource-efficient choice of $k$ is sufficient to enhance existing class-prototype methods when integrated with LayUP.
Method CIFAR IN-R IN-A CUB OB VTAB Cars
ADAM 87.6 72.3 50.5 87.1 74.3 84.3 51.3
ADAM w/ LayUP ($k=4$) 91.0 81.8 63.3 87.5 78.1 93.9
ADAM w/ LayUP ($k=6$) 81.7 82.2
ADAM w/ LayUP ($k=12$) 91.3 87.5 78.1 94.6 81.7
(a) LayUP + ADAM (Zhou et al., 2023b)
Method CIFAR IN-R IN-A CUB OB VTAB Cars
RanPAC 90.7 78.0 58.2 88.5 76.9 92.6 67.5
RanPAC w/ LayUP ($k=4$) 82.4 61.5 88.7 94.0 81.9
RanPAC w/ LayUP ($k=6$) 63.4 88.6 78.3 93.4
RanPAC w/ LayUP ($k=12$) 82.4 78.3 81.7
(b) LayUP + RanPAC (McDonnell et al., 2023)
Table 5: Combination of LayUP with other class-prototype methods. The reported average accuracy scores (%) pertain to the CIL setting as detailed in Sec. 5.1. All baselines were trained using ViT-B/16-IN1K as the pre-trained model and AdaptFormer as the PETL method. Results for RanPAC and ADAM, which were not reported for this specific configuration, have been reproduced using their officially released code repositories. Note that the results for RanPAC and ADAM may differ slightly from those presented in Tab. 1, as the latter reports the best among all configurations of PETL methods and ViT models.
@@ 6 Conclusion
In this paper, we propose LayUP, a simple yet effective rehearsal-free class-prototype method for continual learning with pre-trained models. LayUP leverages multi-layer representations of a pre-trained feature extractor to increase robustness and generalizability, especially under large domain gaps and in low-data regimes. It further computes second-order feature statistics to decorrelate class prototypes, combined with parameter-efficient first session adaptation. Extensive experiments across a range of image classification datasets and continual learning settings demonstrate that LayUP performs strongly against competitive baselines while requiring significantly less memory and computational complexity. Beyond that, our method can be used as a versatile plug-in to improve existing class-prototype methods. In future work, we will further investigate integrating class-prototype methods with continual adaptation of pre-trained models beyond first session adaptation, which is particularly beneficial for learning under multiple distributional shifts. Multi-layer representations serve as a potent source of knowledge that is inherently available in pre-trained models. Based on the findings made in this work, we encourage further exploration into making explicit use of these representations (i.e., reading between the layers) in future approaches to continual learning.